OpenStudy (anonymous):

Given the function f(x) = 2(x + 8), find x if f(x) = 12 2 40 −2 14

OpenStudy (anonymous):

Do you want to see how or do you want me to give you the answer?

OpenStudy (anonymous):

how

OpenStudy (anonymous):

ok so, what times 2 = 12?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

ok then you need to figure out what + 8 = 6

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

12=2(x+8) 2(x+8)=12 x+8=12/2 x+8=6 x=6-8 x=-2

OpenStudy (anonymous):

if it does then that is your answer

OpenStudy (anonymous):

yeah thanks britanny

OpenStudy (anonymous):

yw